OpenStudy (anonymous):

Which line is parallel to the line that passes through the points (2, –5) and (–4, 1)?

Parth (parthkohli):

See any line which has the same slope :)

Parth (parthkohli):

Because parallel line have the same slope :D

Parth (parthkohli):

lines*

OpenStudy (anonymous):

you can use the equation \[y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x - x_1)\] therefore you get \[y - (-5) = \frac{1 - (-5)}{-4 - 2} (x - 2)\] if you simplify.... \[y + 5 = \frac{6}{-6} (x - 2)\] \[y +5 = -(x +2)\] \[y + 5 = -x - 2\] \[y = -x - 7\]